OpenStudy (anonymous):

Put in vertex form and identify vertex h(x)=5x^2-20x+9

OpenStudy (anonymous):

@jim_thompson5910 @Directrix want to take over, sorry i have mid-terms tomorrow (need to study)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

@ghazi @Cha1234 @UnkleRhaukus

jimthompson5910 (jim_thompson5910):

h(x)=5x^2-20x+9 h(x) - 9=5x^2-20x h(x) - 9=5(x^2-4x) h(x) - 9=5(x^2-4x + 4 - 4) h(x) - 9=5(x^2-4x + 4) - 5*4 h(x) - 9=5(x^2-4x + 4) - 20 h(x) - 9=5(x-2)^2 - 20 h(x)=5(x-2)^2 - 20+9 h(x)=5(x-2)^2 - 11 The equation is now in vertex form f(x) = a(x-h)^2+k where a = 5 h = 2 k = -11 so the vertex is (2, -11)

OpenStudy (anonymous):

thank you!!!

jimthompson5910 (jim_thompson5910):

np